FIFTEEN YEAR HENCE, A MAN WILL BE JUST FOUR TIMES AS OLD AS HE WAS 15 YEARS AGO.HIS PRESENT AGE IS?

Answer by checkley79(3341)

x+15=4(x-15)
x+15=4x-60
x-4x=-60-15
-3x=-75
x=-75/-3
x=25 years old is his present age.
Proof:
25+15=4(25-15)
40=4*10
40=40